A parabola passes through the points (7,1), (12,15), (14,9), and (20,t). What is the value of t?
Answer by Alan3354(69443)

A parabola passes through the points (7,1), (12,15), (14,9), and (20,t). What is the value of t?
---------
Find the equation of the parabola.
y = ax^2 + bx + c
Solve for a, b & c using the 3 points given.
49a + 7b + c = 1
144a+12b + c = 15
196a+14b + c = 9
-----------------------
Solve for a, b & c
I would use a matrix:
|y     x^2    x   1|
|y1   x1^2   x1   1|
|y2   x2^2   x2   1| = 0
|y3   x3^2   x3   1|

----
Sub 20 for x in the equation to find t.